OpenStudy (anonymous):

This exercise uses the radioactive decay model. If 250 mg of a radioactive element decays to 180 mg in 60 hours, find the half-life of the element. (Round your answer to the nearest whole number.)

OpenStudy (anonymous):

Formula: Nt = N0 [(1 / 2) ^ (t / t(1/2) ) where Nt = current amount of radioactive substances. N0 = Initial amount of radioactive substances. t1/2 = half life of radioactive substances. t = quantity of substances remains after the time t. in this problem Nt = 189 mg, N0 = 250 mg, , t = 60 hours 200 = 250 [ (1/2)^ (48 / t(1/2) ) ] (1/2)^ (48 / t(1/2) ) = 200/250 = 0.8 48 / t(1/2) = Log ( with the base of 1/2 ) 0.8 = log 0.8 / log 0.5 = 0.32 t(1/2) = 48 /0.32 = 150 so the half life of the radioactive is 150 hours Source: http://answers.yahoo.com/

OpenStudy (anonymous):

@hannahmarie620

OpenStudy (anonymous):

that isn't right